: Alcohol Dehydration Product : Unknown Sample ID Sample Type Injection Volume ISTD Amount :C:\GCsolution Data dehydration reAssign the peaks to the appropriate Alkene structure by comparing the retention time to those of the pure Alkene chromatogram (the one to compare to is the first one that has the structures drawn in pen) .. calculate Peak area for the second page do it for all the peaks except solvent peak which is pentane. Peak area= height*Width@1/2height

Also, Calculate the percent of each compound %X= AreaX/ AreaX+ AreaY+AreaZ*100
Answer #1

ANSWER:

1) Comparing retention time

  • The solvent peak (pentane) is the peak at 0.500 min.
  • The 3-methylcyclohexene appears at 1.322 min
  • The peak at 2.261 min is the 2-methylcyclohexanol
  • Only with the retention times it is difficult identify if 1-methylcyclohexene is the peak at 1.41.3 min or the peak at 1.528 min

2) Calculate peak Area and % of compound

Peak Retention time Height (intensity) Width (min) Area % compound
1 0.344 685 0.151 685 x 0.151 = 103.435 103.435 x 100 / 2516.547 = 4.11 %
2 0.500 (solvent) ---- ---- ---- ----
3 1.322 4398 0.176 4398 x 0.176 = 774.048 774.048 x 100 / 2516.547 = 30.76 %
4 1.413 3954 0.085 3954 x 0.085 = 336.09 336.09 x 100 / 2516.547 = 13.36 %
5 1.528 3533 0.273 3533 x 0.273 = 964.509 964.509 x 100 / 2516.547 = 38.33 %
6 2.261 2435 0.139 2435 x 0.139 = 338.465 338.465 x 100 / 2516.547 = 13.45 %
Total area: 2516.547